### v1

# Learning Interior Point Method for AC and DC Optimal Power Flow

Abstract
This paper proposes a feasibility-guaranteed learning interior point method (L-IPM) to solve both AC and DC optimal power flow (OPF) problems. Given the criticality of OPF, the proposed L-IPM uses a hybrid learning-model approach rather than relying solely on a simple black-box prediction. The traditional IPM follows a central path from an initial point to the optimal solution. However, each iteration involves solving large linear systems, which becomes increasingly expensive as the matrices grow more ill-conditioned in later steps. To address this, we model the IPM trajectory as a time series and train a Long Short-Term Memory (LSTM) network to project the IPM central path using only the first few stable iterations, which carry the most informative features about the path to optimality. We introduce a grid-informed methodology that enforces operational constraints on generation, voltage magnitudes, and line flows to ensure feasibility. The grid-informed LSTM serves as a tool for the IPM central path projection and, followed by a final IPM refinement step, significantly reduces the total number of iterations and time required for convergence. We use a sampling method to generate a wide range of load scenarios to improve generalization across diverse operating conditions, efficiently covering the power system’s operational space. Simulation results on a 2869-bus European high-voltage transmission system show that the proposed L-IPM significantly reduces solution time by up to 94%, while maintaining accuracy and feasibility of the solution. By leveraging early iterations and bypassing the final ill-conditioned and computationally demanding steps of traditional IPM, the proposed L-IPM reduces the number of required iterations by up to 85.5%. Since solution feasibility is also guaranteed, L-IPM outperforms the conventional IPM in both computational efficiency and robustness.
Index Terms:
Learning interior point method, grid-informed LSTM, AC and DC optimal power flow
Nomenclature
Indices and Sets:
$i,j$
Indices for buses
$d$
Indices for loads
$t$
Iteration step
$\mathbb{B}$
Set of all buses
$\mathbb{G}$
Set of generator buses
$\mathbb{L}$
Set of transmission lines
$\mathbb{D}$
Set of loads
System Parameters:
$Y_{ij}$
Entry of the bus admittance matrix
$F_{ij}^{\max}$
Maximum loading of line $(i,j)$
$P_{G_{i}}^{\min},P_{G_{i}}^{\max}$
Generator power limits at bus $i$
$V_{i}^{\min},V_{i}^{\max}$
Voltage magnitude limits at bus $i$
$a_{i},b_{i},c_{i}$
Cost coefficients for generator at bus $i$
Optimization Variables:
$p_{i},q_{i}$
Active/reactive power generation at bus $i$
$v_{i}$
Complex voltage at bus $i$
$\theta_{i}$
Voltage angle at bus $i$
$s_{i}$
Complex power injection at bus $i$
$f_{ij}$
Apparent power flow on line $(i,j)$
Slack, Barrier, and Dual Variables:
$s_{i}^{+},s_{i}^{-}$
Slack variables for generator limits
$v_{i}^{+},v_{i}^{-}$
Slack variables for voltage limits
$\sigma_{ij}$
Slack variable for transmission line limits
$\mu$
Log-barrier parameter in IPM
$\lambda_{i}^{S}$
Multiplier for power balance at bus $i$
$\eta_{i}^{+},\eta_{i}^{-}$
Multipliers for generation limits
$\zeta_{i}^{+},\zeta_{i}^{-}$
Multipliers for voltage constraints
$\xi_{ij}$
Multiplier for thermal line limit on $(i,j)$
GI-LSTM Variables:
$x_{t}$
System state at iteration $t$
$y_{t}$
Predicted OPF solution at time $t$
$C_{t}$
LSTM cell state
$h_{t}$
LSTM hidden state
$f_{t},i_{t},o_{t}$
Forget, input, and output gates in LSTM
$k_{\text{max}}$
Number of IPM iterations used in training
$\theta$
Trainable parameters of the LSTM
$y_{t}^{\text{LSTM}}$
Predicted solution from GI-LSTM
$y_{t}^{\text{final}}$
Final OPF solution after refinement
Loss Function Terms:
$L_{ACOPF}$
grid-informed loss function for ACOPF
$L_{DCOPF}$
grid-informed loss function for DCOPF
$\omega_{1},\omega_{2},\omega_{3}$
Weights balancing prediction accuracy and constraint violation
Data Generation and Sampling:
$P_{d}$
Active load at bus $i\in\mathbb{D}$
$P_{d}^{\text{nominal}}$
Nominal load at bus $i\in\mathbb{D}$
$\alpha_{d}$
Load scaling factor
$n_{d}$
Number of load buses
$n_{\text{total}}$
Total number of load combinations
$n_{\text{samples}}$
Number of Latin Hypercube samples
$X_{i}^{\text{OPF}}$
OPF solution for sample $i$
$\bar{X}^{\text{OPF}}$
Mean OPF solution over samples
$\sigma_{X}^{2}$
Mean squared deviation in OPF results
$Z$
Normal distribution critical value
$\epsilon$
Tolerated estimation error
$\sigma$
Standard deviation
I Introduction
Effective management of economic dispatch, voltage regulation, and system security is key to reliable grid operation, which makes optimal power flow (OPF) an essential tool in modern power systems [1, 2]. As power systems become more complex, the need for fast and scalable OPF solutions becomes more pressing. With the increasing reliance on renewable energy sources, combined with the unpredictable nature of loads and distributed generation, real-time OPF becomes even more challenging to compute. To keep up with these demands, it is necessary to have OPF solvers that are computationally efficient and practical for real-world power systems [3]. These solvers should surpass existing methods in both speed and scalability.
Several methods have been proposed to accelerate OPF computation and improve its efficiency, ranging from advanced mathematical formulations to emerging paradigms such as quantum computing [4, 5, 6]. Among recent algorithmic advancements, techniques such as semidefinite programming [7], quantum-enhanced optimization [8], sequential quadratic programming [9], and methods based on the augmented Lagrangian frameworks [10] have shown considerable promise. In addition to these optimization-based approaches, some researchers have explored hardware and parallelization enhancements, such as GPU acceleration [11] and high-performance computing [12], which can improve performance but require access to specialized infrastructure. Others have used heuristic and metaheuristic algorithms [13], which may offer faster solutions but do not guarantee global optimality.
More recently, attention has turned toward machine learning-based methods, including grid-informed neural networks, supervised learning approximations, reinforcement learning, and graph neural networks [14]. These approaches are typically reduce problem dimensionality or approximate OPF solutions in a black-box manner, often without modifying the internal structure of the OPF or optimization algorithm. While promising, these methods do not always guarantee consistent feasibility across varying system conditions.
Among all OPF solution methods, the interior point method (IPM) remains one of the most popular and reliable techniques [15]. IPM operates by iteratively applying Newton’s steps to optimize system variables, where each step involves solving a large linear system. Typically, the most computationally expensive part of the algorithm is the solution of these linear systems [16]. Some researchers have explored quantum linear solvers to address this computational burden, which offer potential asymptotic speedups [17]. However, hardware limits, errors, and scalability challenges still prevent their use in large-scale OPF [18]. Other approaches to enhance IPM include heuristic methods such as inexact Newton strategies [19] and machine learning-based techniques [14]. Machine learning has primarily been used to provide warm starts for IPM by predicting good initial guesses. While helpful in some cases, these methods do not modify the IPM itself, and as a result, even high-quality warm starts do not consistently lead to reduced IPM iterations and computation time. We want to use machine learning directly to enhance the IPM algorithm’s internal behavior or to learn from its iterative patterns. If the number of Newton steps required for convergence could be reduced through such learning, the overall computational complexity of IPM could be enhanced.
IPM follows a trajectory toward the optimal solution, which can be interpreted as a time series, as it progresses iteratively toward convergence. In this work, we propose to model the trajectory of the central path as a data sequence and learn its behavior using a Long Short-Term Memory (LSTM) network. A key observation motivating this approach is that the most informative patterns in the IPM trajectory typically emerge during the early iterations, while the final iterations primarily serve to refine feasibility and enforce constraint satisfaction. By leveraging this property, we can apply path sampling techniques to project the optimal solution, enabling a significant reduction in the number of Newton steps required and, consequently, a substantial decrease in the overall OPF computation time.
Applying LSTM to project the central path of IPM presents several challenges. Since OPF is a constrained optimization problem, direct projections do not inherently guarantee feasibility. Infeasible solutions can lead to constraint violations, which compromise system reliability and may have catastrophic consequences in real-world power grid operations. To address this, we propose a grid-informed LSTM (GI-LSTM) that incorporates system constraints into the learning process. Another key challenge lies in generating a sufficiently diverse and representative training dataset. For a given power network, the number of possible operating conditions grows exponentially with the number of buses and load variations. Ensuring the generalization and reliability of the learning IPM (L-IPM) approach, therefore, requires careful attention to both data generation and model architecture.
The contributions of this paper are summarized as follows.
  • A novel L-IPM framework is proposed for solving the OPF problem, designed to address these computational and modeling challenges. Our analysis reveals that the early IPM iteration typically corresponds to a system with the lowest condition number. In contrast, subsequent iterations become increasingly ill-conditioned, leading to low accuracy of the solution. Motivated by this insight, we use only the initial few iterations as inputs to the GI-LSTM model.
  • To ensure feasibility, we incorporate a grid-informed loss function that penalizes violations of generator limits, voltage magnitudes, and power flow constraints.
  • To efficiently capture a wide range of system conditions, we employ Latin hypercube sampling (LHS) to generate diverse load scenarios without the need for exhaustive enumeration [20].
  • The GI-LSTM projection is refined using the classic IPM to ensure convergence to an exact and feasible solution.
Simulation results on various systems, including a realistic large 2869-bus system, show that the proposed L-IPM significantly reduces the number of IPM iterations, Newton steps, and solution time while maintaining both solution accuracy and OPF feasibility.
II Preliminaries
This section outlines the standard approach for solving the OPF problem, focusing on the widely adopted formulation that uses IPM as its core optimization technique.
II-A OPF Formulation
OPF minimizes total generation costs subject to power balance and system constraints [21]:
$\min_{\mathbf{p}}\sum_{i\in\mathbb{G}}\left(a_{i}p_{i}^{2}+b_{i}p_{i}+c_{i}\right)$ (1)
where $p_{i}$ is the real power generation at bus $i$, for all $i\in\mathbb{G}$, and $a_{i},b_{i},c_{i}$ are cost coefficients. OPF is subject to AC power flow equations that enforce active and reactive power balance at each bus:
$s_{i}=p_{i}+jq_{i}=v_{i}\sum_{j\in\mathbb{B}}Y_{ij}^{*}v_{j}^{*}$ (2)
where $p_{i}$ and $q_{i}$ are the real and reactive power injections at bus $i$, $v_{i}$ is the voltage at bus $i$, and $v_{j}^{*}$ is the complex conjugate of the voltage at bus $j$. The problem is further constrained by generator limits in (3), voltage constraints 4, and transmission line thermal limits in (5).
$P_{i}^{\min}\leq p_{i}\leq P_{i}^{\max},\quad\forall i\in\mathbb{G}$ (3)
$V_{i}^{\min}\leq v_{i}\leq V_{i}^{\max},\quad\forall i\in\mathbb{B}$ (4)
$|f_{ij}|\leq F_{ij}^{\max},\quad\forall(i,j)\in\mathbb{L}$ (5)
$F_{ij}$ is the apparent power flow on line $(i,j)$.
II-B Classic IPM
IPM, as the state-of-the-art method to solve OPF, introduces slack variables to convert inequality constraints 3 to 5 into equality constraints. It then constructs a Lagrangian function by incorporating all constraints through dual variables. The method adds a logarithmic barrier term to the objective function to ensure non-negativity. IPM solves the resulting system iteratively using the Newton-Raphson method, where each iteration involves solving a system of linear equations to approach the optimal solution. The Lagrangian function $\mathcal{L}$ is formed by incorporating the equality constraints (power balance) using Lagrange multipliers $\lambda_{i}^{P}$ and $\lambda_{i}^{Q}$, and the inequality constraints (generator and voltage limits, thermal limits) using slack variables and their corresponding dual variables. The Lagrangian is given by:
$\mathcal{L}$ $=\sum_{i\in\mathbb{G}}\left(a_{i}p_{i}^{2}+b_{i}p_{i}+c_{i}\right)$
$+\sum_{i\in\mathbb{B}}Re\left\{\lambda_{i}^{S*}\left(s_{G_{i}}-s_{D_{i}}-v_{i}\sum_{j\in\mathbb{B}}Y_{ij}^{*}v_{j}^{*}\right)\right\}$
$+\sum_{i\in\mathbb{G}}\eta_{i}^{+}(p_{i}+s_{i}^{+}-p_{i}^{\max})+\sum_{i\in\mathbb{G}}\eta_{i}^{-}(p_{i}-s_{i}^{-}-p_{i}^{\min})$
$+\sum_{i\in\mathbb{B}}\zeta_{i}^{+}(v_{i}+v_{i}^{+}-V_{i}^{\max})+\sum_{i\in\mathbb{B}}\zeta_{i}^{-}(v_{i}-v_{i}^{-}-V_{i}^{\min})$
$+\sum_{(i,j)\in\mathbb{B}}\xi_{ij}\left(|f_{ij}|^{2}+\sigma_{ij}-(F_{ij}^{\max})^{2}\right)-\mu\sum_{(i,j)\in\mathbb{B}}\log\sigma_{ij}$
$-\mu\sum_{i\in\mathbb{G}}\left(\log s_{i}^{+}+\log s_{i}^{-}\right)-\mu\sum_{i\in\mathbb{B}}\left(\log v_{i}^{+}+\log v_{i}^{-}\right)$ (6)
The term $\Re$ ensures the Lagrangian remains real-valued. To solve the Lagrangian formulation, we construct the Karush-Kuhn-Tucker (KKT) conditions, which are derived with respect to all primal variables (e.g., $\mathcal{PD}=\left\{p_{i},\ q_{i},\ v_{i},\ \theta_{i}\right\}$), slack variables (e.g., $\mathcal{SD}=\left\{s_{i}^{\pm},\ v_{i}^{\pm},\ \sigma_{ij}\right\}$), and dual variables associated with the constraints. Newton’s method is then applied to solve this system iteratively. At each iteration, a linear system is formed with a coefficient matrix known as the KKT matrix.
III Proposed Grid-Informed L-IPM
The IPM-based OPF follows a structured trajectory known as the central path, illustrated in Fig. 1. We have observed that the initial iterations of the central path contain valuable information about its trajectory and, consequently, the final optimization solution. These early steps typically cover a significant portion of the path toward optimality. One can also clearly recognize that the IPM central path closely resembles a time series. Given this characteristic, LSTM networks are well-suited for leveraging early iteration data to predict the solution.
Furthermore, as Fig.2(a), early iterations tend to involve matrices with relatively low condition numbers, making them computationally more stable and efficient to solve. In contrast, the final iterations often involve ill-conditioned matrices, which are more numerically sensitive and generally more expensive to solve. We observed that in many cases, the time needed to solve the IPM iterations increases toward the final iterations, as shown in Fig. 2(b). However, a higher condition number does not necessarily lead to longer solution times, as the computational effort also depends on factors such as matrix sparsity, fill-in during factorization, and solver efficiency [12]. This behavior can also be attributed to the efficient sparse solvers and numerical techniques implemented in MATPOWER.
By leveraging early iterations and avoiding direct computation of poorly conditioned steps, we propose an L-IPM approach that significantly reduces the total IPM time required to obtain the OPF solution. The proposed L-IPM framework is illustrated in Fig. 3.
Figure 1: IPM path to the optimal solution.
((a))
((b))
Figure 2: Condition number and solution time of IPM’s iterations in DCOPF.
Figure 3: L-IPM framework.
III-A Learning Model Architecture
An LSTM network can be trained to mimic the IPM central path trajectory by learning from the initial iterations obtained through solving the Newton step. Such an LSTM network consists of multiple layers that extract temporal features from past system states, followed by a fully connected layer that maps these features to OPF variables [22].
At each time step $t$, the LSTM cell state updates based on previous iterations:
$C_{t}=f_{t}\cdot C_{t-1}+i_{t}\cdot\tilde{C_{t}}$ (7)
$h_{t}=o_{t}\cdot\tanh(C_{t})$ (8)
where $f_{t}$, $i_{t}$, and $o_{t}$ represent the forget, input, and output gates, respectively, controlling the flow of information within the network. The cell state $C_{t}$ stores long-term dependencies, while the hidden state $h_{t}$ provides a compressed representation for predicting OPF solutions. For each time step $t$ in (7), we solve OPF using MATPOWER and collect a dataset as structured in Table I.
TABLE I: Dataset Structure for GI-LSTM Training
T Step Bus ID Load (MW) Gen (MW) Volt. (p.u.) Ang. (°) Feas.
1 1 100.5 120.3 1.02 3.5 Yes
2 1 101.8 119.7 1.01 3.6 Yes
$\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$
III-B LSTM Training Mapping Strategy
Given a set of system conditions, the LSTM model is trained to learn a temporal mapping function from a sequence of past states to the final OPF solution [22]. Formally, the model learns:
$y_{t}=f_{\theta}(x_{t-k},x_{t-k+1},...,x_{t})$ (9)
where $x_{t}$ represents the system state at time $t$, including features such as load demand, generation constraints, and historical OPF results, and $y_{t}$ denotes the predicted solution. The model parameters $\theta$ are optimized to minimize the prediction error while encouraging feasible outputs.
In the context of LSTM, this mapping is further specialized to learn the relationship between early-stage iterations of IPM and its final converged solution. Let $k_{\text{max}}^{0}$ denote the number of early IPM iterations used as input, which in our case is 3 iterations. Then, the input-output mapping becomes:
$X=\{x_{1},x_{2},\dots,x_{k_{\text{max}}^{0}}\},\quad Y=x_{T}$ (10)
Each $x_{k}$ represents the system state at iteration $k$, including generator outputs, bus voltages, Lagrange multipliers, and slack variables. The LSTM model learns a function $f_{\theta}(X)=Y$, where the parameters $\theta$ are defined as the model’s learnable parameters.
III-C Grid-Informed LSTM for OPF Feasibility
The LSTM network projects the final iteration of the IPM, which is expected to lie close to the optimal OPF solution. This is illustrated in Fig. 4, which provides a zoomed-in view of the optimal solution shown in Fig. 2(a). While LSTM is effective in learning the convergence trajectory of IPM, ensuring that the LSTM projection of the IPM central path remains feasible or stays within the green region surrounding the optimal solution is critical for maintaining OPF feasibility and satisfying power system operational constraints. If the projection falls within the red zone, it would lie outside the OPF feasible region, thus violating the operational constraints. This is problematic because, as discussed later, the proposed L-IPM approach relies on a hybrid learning-model framework that uses the learner’s projection to guide IPM refinement near the optimal solution. If the learner’s projection is infeasible and falls outside the feasible region, the IPM may struggle to refine it and converge to the true OPF optimal solution.
Several key challenges must be addressed to enhance the robustness of L-IPM, including capturing the nonlinear convergence behavior of IPM, maintaining predictive accuracy, and ensuring generalization across diverse operating conditions. We address these challenges using the proposed IPM-tailored GI-LSTM framework.
Figure 4: GI-LSTM projection of IPM central path.
A standard mean-squared error loss optimizes prediction accuracy but does not ensure physically feasible OPF solutions. To enforce feasibility, a grid-informed loss function for ACOPF and DCOPF is introduced in (11) and (12), respectively:
$L_{ACOPF}=$ $\ \omega_{1}\sum_{i\in\mathbb{G}}\left(P_{G,i}^{\text{LSTM}}-P_{G,i}^{\text{IPM}}\right)^{2}+$ (11)
$\ \omega_{2}\sum_{i\in\mathbb{B}}\left(V_{i}^{\text{LSTM}}-V_{i}^{\text{IPM}}\right)^{2}+$
$\ \omega_{3}\sum_{(i,j)\in\mathbb{L}}\max\left(0,F_{ij}^{\text{LSTM}}-F_{ij}^{\max}\right)$
$L_{DCOPF}=$ $\ \omega_{1}\sum_{i\in\mathbb{G}}\left(P_{G,i}^{\text{LSTM}}-P_{G,i}^{\text{IPM}}\right)^{2}+$ (12)
$\ \omega_{2}\sum_{i\in\mathbb{B}}\left(\theta_{i}^{\text{LSTM}}-\theta_{i}^{\text{IPM}}\right)^{2}+$
$\ \omega_{3}\sum_{(i,j)\in\mathbb{L}}\max\left(0,P_{ij}^{\text{LSTM}}-P_{ij}^{\max}\right)$
where the first two terms ensure accurate predictions for generator values and bus voltage magnitudes, while the third penalizes constraint violations in power flows. The hyperparameter $\omega$ balances accuracy and feasibility. The first two terms $\omega_{1}$ and $\omega_{2}$ ensure accurate predictions of generator outputs and voltage magnitudes, while the third penalizes power flow constraint violations. Among these, $\omega_{3}$ is typically assigned the highest value to strongly discourage infeasible solutions.
III-D GI-LSTM Tuning
III-D1 Number of GI-LSTM Layers
A two-layer GI-LSTM architecture effectively captures IPM iteration sequences for small- to medium-scale OPF problems and larger networks under typical conditions. For large-scale OPF cases with more complex IPM convergence trajectories, a three-layer GI-LSTM is recommended to better model these dynamics. This configuration balances modeling depth and computational efficiency.
III-D2 Regularization Techniques
To prevent overfitting and enhance model generalization across diverse OPF scenarios, regularization strategies are used. Dropout regularization randomly deactivates neurons during training, with an appropriate dropout rate selected to balance robustness and information retention. Additionally, L2 regularization (weight decay) penalizes large weights, simplifying the model to improve generalization [22].
III-E L-IPM for OPF Solution
Reducing the number of IPM iterations before predicting the final solution allows for substantial computational savings. However, it is important to ensure that the GI-LSTM model generalizes well to diverse OPF conditions. Using L-IPM, which combines GI-LSTM predictions with IPM refinement, we ensure the feasibility and accuracy of the final solution. The process begins by training GI-LSTM on historical OPF solutions obtained from IPM. Once trained, GI-LSTM generates an initial guess $y_{t}^{LSTM}$ for the OPF solution, which is then refined using a final IPM optimization step:
[…]
The proposed L-IPM framework modifies the classic IPM algorithm to achieve faster convergence toward the same solution. Feasibility and accuracy are preserved through a final IPM refinement step, as illustrated in Fig. 3.
IV Data Generation Techniques
The L-IPM model’s effectiveness depends on the training dataset’s diversity and representativeness, particularly in capturing the impact of load variations on the power system. The key objective is to determine the minimum number of load scenarios required to ensure generalization while avoiding excessive computational burden. We suggest a systematic approach to estimate the required number of load cases based on power system size, load variability, and OPF sensitivity.
IV-A Defining Load Scenarios and Their Variability
Let $\mathbb{D}\subseteq\mathbb{B}$ denote the subset of load buses where active and reactive power demand varies. The power system exhibits different operating conditions due to changes in load levels. These variations are captured using a probabilistic model:
$P_{d}=P_{d}^{\text{nominal}}\cdot\alpha_{d},\quad\forall d\in\mathbb{D}$ (14)
where $P_{d}$ is the active power demand and $P_{d}^{\text{nominal}}$ is the nominal demand at bus $d$. The scaling factor $\alpha_{d}$ is modeled as a random variable drawn from a uniform distribution $U(a,b)$, where the bounds $(a,b)$ depend on the considered load scenario. Specifically, $\alpha_{i}\sim U(0.9,1.1)$ for the nominal load condition, $\alpha_{i}\sim U(1.1,1.2)$ under the high-demand scenario, $\alpha_{i}\sim U(0.8,0.9)$ for the low-demand scenario, and $\alpha_{i}\sim U(0.8,1.2)$ for the mixed load scenario.
Each bus can assume one of the predefined load scaling factors for a system with $n_{d}$ load buses. A fully exhaustive enumeration of all possible load combinations results in:
$n_{\text{total}}=m^{n_{d}}$ (15)
where $m$ is the number of discrete scaling values between 0.8 and 1.2. If continuous variation is considered, the number of scenarios would be infinite. Hence, due to the exponential growth in scenarios, we need to use a method that realistically covers all possible load scenarios.
To ensure $\alpha_{d}$ varies across all load buses, we adopt LHS, a stratified sampling method that evenly partitions the input space and selects representative samples from each interval. Unlike random sampling, which may cluster points unevenly, LHS provides diverse and well-distributed load conditions [20].
IV-B Minimum Required Load Scenarios for Different System Sizes
The minimum required number of load cases for even coverage in LHS is determined using the Monte Carlo error bound approach or the central limit theorem [23]:
$n_{\text{samples}}\geq\left(\frac{Z\sigma}{\epsilon}\right)^{2}$ (16)
where $Z$ is the critical value from the normal distribution for a 95% confidence level, $\sigma$ is the assumed normalized standard deviation ($\approx 1$), and $\epsilon$ is the desired relative error tolerance.For a more conservative estimate in large systems, a common heuristic suggests setting the total number of samples as $n_{\text{samples}}=10\times n_{d}$.
The variance in OPF solutions is analyzed as a function of the number of load cases to validate that the generated scenarios are sufficient. If increasing $n_{\text{samples}}$ beyond a certain point does not significantly change the OPF results, then additional cases are unnecessary. The mean squared deviation between OPF solutions is shown as:
[…]
where $\bar{X}^{\text{OPF}}$ is the mean OPF solution across samples.
  • If $\sigma_{X}^{2}$ stabilizes with increasing $N_{\text{samples}}$, the dataset is considered sufficient.
  • If $\sigma_{X}^{2}$ continues to decrease, more samples are required.
Using these guidelines, Table II presents the estimated number of load scenarios. These values ensure robust GI-LSTM generalization across system sizes, improving both accuracy and computational efficiency.
TABLE II: Estimated Number of Load Scenarios
System Size $n_{d}$ $n_{\text{samples}}$
Small $n_{d}\leq 50$ 1000
Medium $50\leq n_{d}\leq 500$ 1000–2000
Large $n_{d}>500$ 2000–10000
V Numerical Results
V-A Simulation Setting
To evaluate the effectiveness and scalability of the proposed L-IPM, we conducted tests on power systems of varying sizes. Our analysis began with a small-scale 3-bus system and a standard 24-bus system, followed by larger systems comprising 118 and 2869 buses to assess the scalability of the method. The GI-LSTM model was implemented in Python, while the IPM-based OPF simulations were carried out using the MATPOWER 8.0 toolbox in MATLAB.
For the performance evaluation of the GI-LSTM model, we used multiple commonly used metrics, including mean squared error (MSE), mean absolute error (MAE), root mean squared error (RMSE), $R^{2}$ score, loss function trends, and residual analysis. We also compared the performance of the classic OPF and the L-IPM-based OPF using metrics, including the objective function value, feasibility condition error, complementarity condition error, and objective function error. These evaluations were conducted under various load scenarios, including adverse conditions with multiple active inequality constraints.
V-B Look-Back Window Determination
To determine the appropriate number of early iterations to input into the GI-LSTM, we conducted multiple simulations. A single iteration does not provide enough historical context for the GI-LSTM to capture temporal patterns and interdependencies. Therefore, we increased the number of early iterations, or the look-back window, from one to five and observed that in many cases, the output became more accurate with a higher number of iterations. However, there were instances where increasing the number of iterations led to a slight decrease in accuracy. In Fig. 5, we compare the $R^{2}$ scores for two systems. We observe that setting the look-back to three provides satisfactory performance, and for the 2869-bus system, it achieves the best performance overall. Moreover, in many cases, increasing the look-back from three to four or five does not significantly improve precision. Based on these observations, we selected a look-back of three iterations to feed into the GI-LSTM.
Figure 5: Comparison of R2 scores for different look-back values.
V-C GI-LSTM Performance
The evaluation metrics of GI-LSTM demonstrate its effectiveness. For instance, the loss function of the trained GI-LSTM on two test systems is shown in Fig. 6. We observed that both the training and validation losses converged toward small values, indicating that the GI-LSTM model was effectively optimized and capable of learning meaningful patterns from the data. A high coefficient of determination $R^{2}$ across all four test systems further confirms the model’s predictive accuracy, as it demonstrates a strong alignment between the predicted and actual values in the validation phase. These high-accuracy projections are refined with IPM to improve numerical stability. The performance metrics for each system are summarized in Table III, highlighting the consistency and generalizability of the proposed approach across different system sizes.
((a))
((b))
Figure 6: GI-LSTM loss function error.
TABLE III: R2 Score for GI-LSTM Across Different Systems
3-bus 24-bus 118-bus 2869-bus
ACOPF 0.981 0.977 0.994 0.902
DCOPF 0.990 0.971 0.983 0.949
In Table III, we observe that the R2 score for ACOPF in the 2869-bus system drops to 0.9. This decline is primarily attributed to the nature of the GI-LSTM model, which attempts to enforce all solutions within the feasible region. However, we later observed that not all demand scenarios have feasible solutions when solved using IPM. Therefore, the observed drop is largely because the solutions remain constrained within the feasible region. Table IV shows additional metrics used to evaluate the trained GI-LSTM model, further supporting our approach’s strong performance and reliability.
TABLE IV: GI-LSTM Prediction Errors Across Different Systems
3-bus 24-bus 118-bus 2869-bus
MSE 0.010 0.012 0.004 0.080
MAE 0.068 0.067 0.033 0.120
RMSE 0.100 0.110 0.068 0.280
V-D GI-LSTM Outpout
The GI-LSTM model predicts outputs that are both accurate and close to the exact solution. For instance, as Fig. 7, the results generated by GI-LSTM closely match those obtained using IPM, which demonstrates the model’s strong predictive capability.
((a))
((b))
Figure 7: GI-LSTM output comparison with IPM in the 3-bus system.
However, in power system operations, ensuring feasibility is critical. Even minor violations of operational limits—such as voltage, power flow, or generation constraints—can pose serious risks to system stability and reliability. To address this, we refine the GI-LSTM predictions by implementing IPM using the GI-LSTM projected central path. This step ensures that the final solution remains almost the same as the optimal OPF solution and strictly satisfies all system physical and operational constraints. Hence, the feasibility of the L-IPM solutions for OPF is guaranteed.
V-E L-IPM
Simulation Results Summary: The key observation from the L-IPM-based OPF evaluation is highlighted below. Table V summarizes the number of iterations required to solve the OPF problem across various load scenarios. We report the minimum and maximum number of iterations observed during the simulations for each test system. As the system size increases, IPM exhibits higher sensitivity to load variations, resulting in a larger gap between the minimum and maximum number of iterations. This behavior is primarily due to more active inequality constraints under stressed conditions. In contrast, such variation is not observed in the number of L-IPM iterations, highlighting its robustness to load fluctuations. The percentage of average iteration reduction in L-IPM compared to classic IPM is shown in Fig 8(c). For instance, in the case of ACOPF on the 2869-bus system, L-IPM reduced the number of iterations by an average of 85.5%, which is a substantial improvement compared to the classical IPM state.
TABLE V: OPF Iteration Comparison using IPM and L-IPM
DCOPF ACOPF Enhancement (%)
min max min max min max
3-bus 10 (4) 10 (4) 11 (4) 12 (4) 60 66.7
24-bus 8 (4) 37 (5) 13 (4) 16 (4) 50 86.5
118-bus 9 (4) 11 (4) 14 (4) 15 (4) 55.6 73.4
2869-bus 11 (4) 34 (5) 25 (7) 55 (8) 63.7 85.5
Values in parentheses indicate the number of L-IPM iterations.
((a))
((b))
((c))
Figure 8: Performance comparison of IPM and L-IPM.
The iterations avoided by the proposed L-IPM correspond to the most time-consuming and computationally challenging phases of the classic IPM. These later iterations are typically associated with ill-conditioned systems that are computationally expensive to solve, as shown in Fig. 2(a). The observed reduction in runtime exceeds the reduction in iteration count. Table VI reports the total runtime for different test cases under worst-case demand scenarios. It compares the performance of IPM and L-IPM for both AC and DCOPF. As seen, L-IPM achieves substantial runtime savings across all test cases. Fig. 8(a) and (b) visualize the computational time comparison between IPM- and L-IPM-based OPF solutions, performed on a machine with 16 GB RAM and an Intel Core i9 processor. Given the different solution time scales, and for improved visualization, we present the results for the 2869-bus system in a separate figure. A significant reduction in runtime is observed across all cases for both ACOPF and DCOPF problems, particularly for larger, real-world-scale systems. For example, in the case of ACOPF on the 2869-bus European high-voltage transmission system in Fig. 8(b), the proposed L-IPM-based approach is 94% faster than the classical IPM-based OPF in finding the same solution as classical IPM.
TABLE VI: OPF Runtime Comparison of IPM and L-IPM
IPM (s) L-IPM (s) Time Reduction (%)
AC DC AC DC AC DC
3-bus 0.53 0.25 0.103 0.029 80.6 88.3
24-bus 0.25 0.084 0.105 0.015 57.8 81.6
118-bus 0.85 0.32 0.08 0.017 90.5 94.5
2869-bus 12.33 1.31 0.73 0.126 94.1 90
Below, we present a detailed analysis of each test system to provide further insights into the simulation results.
V-E1 3-bus system
We report detailed results for three demand scenarios (DSs): low load, medium load, and high load. In Figs. 9(a) and (c), the classic IPM is used to solve the DSs, and the OPF objective function values are reported across the iterations. Figs. 9(b) and (d) show the performance of L-IPM, where the GI-LSTM output, trained based on observing just the first three iterations, is refined using IPM. L-IPM reaches the same solution as IPM but with fewer iterations, highlighting how L-IPM helps accelerate convergence while maintaining accuracy.
((a))
((b))
((c))
((d))
Figure 9: OPF objective function values for the 3-bus system.
For a more detailed comparison, we evaluated the classical IPM and the proposed L-IPM feasibility and complementarity conditions. As Fig. 10, L-IPM significantly reduces the error to below 10e610e-6 after just one IPM iteration. We analyzed the structure of the Newton steps, particularly the $A$ matrix in the linear system $Ax=b$, and observed that most of the significant changes occur during the early iterations. The remaining iterations primarily refine the solution by accurately identifying violated constraints and adjusting the slack variables within the IPM framework. The L-IPM learns the behavior of these variables, and with a carefully selected set of load scenarios, as described in the data generation section, it develops an understanding of the feasible region of the system, including generation limits and line capacities.
Figure 10: Feasibility and complementarity condition comparison between IPM and L-IPM in 3-bus system.
In the 3-bus system case, we also observed that most solution trajectories were similar. The main difference was in the objective function value. This consistent and well-behaved nature of IPM helped L-IPM accurately capture and predict solutions that are very close to the exact optimal values.
V-E2 24-bus system
For this system, we observed that a higher number of IPM iterations was often required to reach a solution. In particular, under extreme load scenarios, the solver needed more than 30 iterations in many cases.
As illustrated in Fig. 11, we compare the performance of ACOPF and DCOPF for two representative cases. The first case involves a load scenario close to the nominal operating point, while the second represents an extreme load condition. The figure illustrates the increased number of iterations required by IPM to solve this load scenario. For both load scenarios, L-IPM converges in just four iterations for ACOPF and five iterations for DCOPF.
((a))
((b))
((c))
((d))
Figure 11: OPF objective function values for the 24-bus system.
To better illustrate how efficiently L-IPM reduces the objective function relative error compared to IPM, Fig. 12 presents the relative error of L-IPM for both DCOPF and ACOPF. Fig. 13 presents a comparison of feasibility and complementarity condition errors under an extreme load scenario, further highlighting the robustness and efficiency of L-IPM.
((a))
((b))
Figure 12: L-IPM OPF objective function error of the 24-bus system.
In extreme load cases, we found that L-IPM needs up to two IPM iterations to converge. This suggests that while GI-LSTM outputs may need slightly more effort to handle highly stressed scenarios, it still provides a significantly faster solution compared to solving the problem entirely with the traditional IPM. Even in these challenging cases, convergence was achieved with only a few iterations.
Figure 13: Feasibility and complementarity condition comparison between IPM and L-IPM in 24-bus system.
V-E3 118-bus system
We selected this system to explore whether larger systems tend to exhibit similar convergence patterns as seen in the 24-bus system. The relatively stable behavior observed in the 118-bus case suggests that convergence characteristics may vary depending on the specific structure and properties of each power system, rather than system size alone.
As Figs. 14(a) and (c) show, ACOPF and DCOPF exhibit stable and well-behaved convergence, even under extreme load scenarios. The system can tolerate significant load variations without any noticeable degradation in performance. The L-IPM also performed reliably, as Fig. 14(b) and (d), where the results for two representative load scenarios demonstrate that the model had no difficulty in finding accurate and feasible solutions.
((a))
((b))
((c))
((d))
Figure 14: OPF objective function values for the 118-bus system.
V-E4 2869-bus system
To evaluate the scalability of our approach, we applied it to a large-scale 2869-bus system. The ACOPF formulation includes 6758 variables, along with 13,256 dual variables corresponding to IPM constraints. Due to the scale of the problem, IPM faces increased difficulty in convergence. While the method typically identifies a viable OPF solution within a few iterations, the subsequent iterations are primarily used to detect the active inequalities and enforce the physical constraints of the system.
As shown in Fig. 15, L-IPM significantly reduces the number of Newton steps required to reach convergence. Given that the dimension of the Newton linear system reaches 20,014 in this system, the computational time needed to solve each Newton step becomes substantial and directly impacts the total OPF solution time. Retaining the feasibility of the power system is critically important, as any infeasible solution can lead to severe operational and economic consequences. As shown in Fig. 16, L-IPM rapidly reduces the feasibility error, demonstrating its effectiveness in guiding the solution toward a feasible and reliable operating point.
((a))
((b))
((c))
((d))
Figure 15: OPF objective function values for the 2869-bus system.
Figure 16: IPM and L-IPM Feasibility and complementarity condition comparison in 2869-bus system.
Given the large scale of the system, load scenarios have a significant impact on the final solution. In particular, IPM struggled to find a feasible solution under poorly conditioned load scenarios, even after many iterations. The solver sometimes reached as many as 53 iterations without convergence. As illustrated in Fig. 17(a), we simulated a challenging load scenario where neither IPM nor L-IPM was able to converge. Even in scenarios where no feasible solution exists, L-IPM can reach this outcome more quickly than the standard IPM. When such a poorly conditioned load scenario is input into our proposed method, the grid-informed structure of GI-LSTM helps ensure that its output remains within or close to the feasible region. After the IPM refinement step, we observed that the IPM fails after approximately five iterations, returning a numerical failure error, similar to the behavior seen in the standalone IPM run. The trend of the feasibility error under this scenario is presented in Fig. 17(b), further highlighting how the hybrid approach accelerates the detection of infeasibility while still adhering to system constraints as closely as possible.
((a))
((b))
Figure 17: IPM and L-IPM comparison in an extreme demand scenario.
VI Conclusion
This paper proposes a feasibility-guaranteed learning interior point method (L-IPM) aiming to accelerate the runtime of AC and DCOPF problems. While early IPM iterations carry valuable information about the optimal solution, the later ones mainly serve to ensure feasibility. Leveraging this observation, we propose an L-IPM that reduces the number of costly iterations. Our method uses a GI-LSTM to project the IPM central path and find a feasible point near the optimal solution. A final IPM refinement stage confirms feasibility and improves solution accuracy.
To evaluate the proposed approach, we tested it on four systems ranging from small (3- and 24-bus) to medium and large (118- and 2869-bus) networks. These case studies demonstrate the scalability, accuracy, and computational efficiency of our method. Results show that L-IPM, when trained on carefully selected load scenarios, significantly accelerates convergence, both in runtime and in the number of IPM iterations. Under challenging load conditions, L-IPM consistently outperformed traditional IPM by requiring fewer iterations. In many cases, it reduced runtime by about 90% compared to the classic IPM. Further analysis of feasibility and complementarity conditions confirmed the robustness of the proposed method, as it rapidly reduced constraint violations while maintaining solution accuracy. Although some extreme load scenarios—particularly in the 2869-bus system—introduced challenges, L-IPM still identified infeasibility more quickly than conventional IPM.
Future work could focus on enhancing the model’s robustness under extreme operating conditions, incorporating component outages, and accounting for the stochastic nature of renewable energy sources and demand forecasting to improve performance under uncertainty.